OpenStudy (kendellfawn):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>?

hartnn (hartnn):

do you know the formula?

OpenStudy (kendellfawn):

No idea, this lesson is all over the place, covers a lot. Mind walking me through this?

hartnn (hartnn):

sure you may use this one, though there are other ways \(u.v = |u||v|\cos \theta \)

OpenStudy (kendellfawn):

Okay, so how would I go about setting this up?

hartnn (hartnn):

you know how to find u.v, right?

hartnn (hartnn):

|u| is the magnitude of vector u

OpenStudy (kendellfawn):

Yes. Okay.

hartnn (hartnn):

so what do u get for |u| |v| and u.v ?

OpenStudy (mathmale):

Look up "vector magnitude" or some similar search term. Given vector u=<6, -1>, how would you find its magnitude? For reference, this magnitude is \[\sqrt{6^2+1^2)}.\] Verify that. Find the magnitude (length) of vector v also.

OpenStudy (mathmale):

hartnn was encouraging you to use the dot product of vectors u and v, and he has given you the proper formula for this. Any questions about "dot product?"